# SLR: Automated Synthesis for Scalable Logical Reasoning

### v4

@@ 4 SLR-Bench: Instantiating SLR
The logic task is inspired by the V-LoL trains domain helff2025vlol; michalski1980; Mitchell1997MachineLI, chosen for three main reasons. First, its hierarchical object structure (trains $\rightarrow$ cars $\rightarrow$ attributes) naturally gives rise to first-order rules that are far richer than simple lookups, yet more tractable than general theorem proving. Second, every attribute has a small, discrete grounding domain, which allows us to measure and control the complexity of the problem precisely. Third, we ensure expandability and novelty, as SLR-Bench is fully synthetic and programmatically generated.
@@
Languages.
Each curriculum level is parameterized by level-specific language $\mathcal{L}$, detailed in App. A.1. The vocabulary includes mutually exclusive class labels eastbound and westbound, which serve as the targets for classification tasks. Additionally, the vocabulary includes five predicates (has_car, car_num, car_color, car_len, and has_wall) with their respective grounding domains specified in App. A.1. As curriculum levels increase in complexity, the vocabulary expands monotonically by introducing new predicates and grounding domains selected from a predefined set. These include categorical predicates such as has_roof, has_payload, has_window, car_type, and numerical predicates load_num, has_wheel, passenger_num. Semantic coherence is ensured by constraining predicate groundings to valid combinations (e.g., only colors as arguments for car_color), and by enforcing mutually exclusive constraints across predicates (e.g., passenger cars cannot carry payloads).
Task configs.
Each curriculum level is parameterized by level-specific settings of $\theta$, summarized in App. Tab.2 and supplied directly to the synthesizer (Alg.1). Problem size ($\kappa$) increases steadily across levels, maintaining an equal balance of positive and negative samples. Levels 1–5 use a mirror sampling policy for background knowledge, generating simple, nearly identical east- and westbound trains that differ only in ground atoms relevant to $R^{\star}$. From level 6 onward, the background is sampled uniformly from the filtered Herbrand base, increasing diversity. Rule generation is uniform for the basic levels; from level 6, 30% of rules are LLM-guided, introducing greater variety in variables, arithmetic, recursion, and more.
⟦t⟧ Curriculum Learning (SLR-Bench) LRL ($\uparrow$0-20) Syntax($\uparrow$%) Basic($\uparrow$%) Easy($\uparrow$%) Medium($\uparrow$%) Hard($\uparrow$%) Llama-8b-it 3.6 99 61 10 1 0 Llama-8b-it-SLR 8.8 (+5.2) 100 (+1) 96 (+35) 56 (+46) 20 (+19) 5 (+5) Downstream Reasoning Performance MMLU ($\uparrow$%) MMLU-Stats ($\uparrow$%) MMLU-CS ($\uparrow$%) MMLU-ML ($\uparrow$%) LogiQA ($\uparrow$%) LogiQA2 ($\uparrow$%) Llama-8b-it 63.3 42.6 61.0 49.1 30.1 34.3 Llama-8b-it-SLR 66.1 (+2.8) 59.7 (+17) 75.0 (+14) 52.7 (+3.6) 31.0 (+0.9) 39.4 (+5.2) GPQA ($\uparrow$%) GPQA-Ext. ($\uparrow$%) GPQA-Dia. ($\uparrow$%) ARC-Easy ($\uparrow$%) ARC ($\uparrow$%) HellaSwag ($\uparrow$%) Llama-8b-it- 31.7 26.9 21.7 81.4 52.7 57.4 Llama-8b-it-SLR 32.8 (+1.1) 33.0 (+6.1) 28.3 (+6.6) 82.8 (+1.4) 54.6 (+1.9) 58.9 (+1.5)
Table 3: Curriculum Learning and Generalization. Benchmark scores ($\uparrow$%) for base and SLR-tuned models (Llama3.1-8b-it) on SLR-Bench and downstream benchmarks; LRL measuring cumulative curriculum progress. The tuned model surpasses the baseline across all curriculum stages, while generalizing to other reasoning tasks.
Curriculum.
SLR-Bench comprises 20 levels across four tiers: basic, easy, medium, and hard. Each level systematically increases complexity by expanding task size ($\kappa$), adding new car constants and predicates, lengthening rules, and varying both the background knowledge and rule sampling policy; see App.Sec.A, Tab. 2. As a result, the combinatorial space of possible tasks grows exponentially, and later levels become progressively harder and require deeper reasoning beyond surface cues.
Intended Use.
SLR-Bench is designed for two complementary purposes. (1) As a static benchmark, it enables fine-grained evaluation of an LLM’s reasoning abilities across tasks of increasing logical complexity. It is also easily extensible to accommodate future improvements in model capabilities. (2) As a dynamic curriculum, it serves as a training backbone, supplying structured reasoning tasks and feedback to enhance reasoning in both conventional and reasoning LLMs.
5 LLMs Can’t Do Induction at Scale
We evaluate and train LLMs on SLR-Bench, assessing reasoning, syntactic correctness, and computational efficiency across four difficulty levels: basic, easy, medium, and hard. Our analysis highlights key trends, common failure modes, and the effectiveness of curriculum-based logic-tuning.
Training Setup.
We investigate how LLMs benefit from curriculum training on SLR-Bench with SFT (for more details see App. Sec. G). To prevent data leakage, we ensure that no prompts or rules from the test set are included in the training set.
Evaluation Setup.
All models are evaluated in a zero-shot setting using SLR-Bench prompts, with a single attempt per task (pass@1). We report the following metrics:
(i) Logical Reasoning Level (LRL): The cumulative model score over all curriculum levels $L$, where $\text{solved}_{\ell}$ and $\text{tasks}_{\ell}$ denote the number of solved and total tasks at level $\ell$, respectively:
$\text{LRL}=\sum_{\ell=1}^{L}\frac{\#\text{solved}_{\ell}}{\#\text{tasks}_{\ell}}$
(ii) Syntax Score: The proportion of predicted logic rules that are syntactically valid.
(iii) Logical-Reasoning Accuracy: The fraction of correct solutions per complexity tier.
(iv) Compute: The aggregate completion tokens and computational cost for each the models.
For further details on downstream evaluations, see App. Sec. G, and for pricing, refer to Tab. 5.
5.1 Analysis and Key Findings
In the following, we highlight downstream gains from training via SLR curriculum learning (Tab. 3). Next, we benchmark SOTA conventional and reasoning LLMs on SLR-Bench (Tab. 4), showcasing the trade-offs of test-time compute (Fig. 5).
SLR Boosts Downstream Reasoning.
Curriculum learning on SLR-Bench yields substantial improvements in both in-domain and downstream reasoning (see Tab 3). SLR-tuned models outperform all conventional LLMs on SLR-Bench (cf. Tab.4) and surpass reasoning LLMs such as Gemini-2.0-flash-thinking, while using far fewer inference tokens and compute resources. On popular reasoning benchmarks, SLR delivers gains on logic-intensive tasks, e.g., on MMLU High School Statistics (+17), Computer science (+14), and Machine learning (+3.6) hendryckstest2021, as well as notable improvements on LogicQA liu2020logiqa, LogicQA2 Hanmeng2023logicqa2, ARC Clark2018ThinkYH, HellaSwag zellers2019hellaswag, GPQA, and GPQA-Extended, GPQA-Diamond rein2024gpqa. These consistent gains across curriculum levels and downstream benchmarks demonstrate that curriculum learning with SLR not only enhances in-domain reasoning but also generalizes effectively to diverse reasoning tasks.
[…]
Curriculum Levels Modulate Task Complexity: LLMs Break Down as Complexity Increases.
SLR-Bench creates a controlled gradient in logical complexity as model performance steadily declines throughout the curriculum levels (cf. Fig. 3, Tab. 4). Most models readily solve the basic levels. However, base LLMs already struggle on the easy tasks, solving less than half. Reasoning LLMs provide improved performance, though they incur substantial drops at the medium levels. Only a few reasoning LLMs manage to solve more than half the medium tasks, yet none on hard. This pattern is also reflected in the LRL score, empirically indicating how far each model can progress before performance collapses.
Reasoning Remains Challenging; Syntax Not.
Base LLMs reliably generate syntactically valid rules, reflected in their high syntax scores (see Tab. 4). Reasoning models exhibit slightly lower scores, particularly on more complex reasoning tasks, where longer outputs can lead to invalid or missing responses. Nonetheless, the primary barrier to higher performance is semantic, as reflected by the gap between syntax and LRL in Tab. 4.
Scaling Test-time Compute Improves Reasoning, but Returns Diminish and Costs Escalate.
Reasoning LLMs clearly outperform the base models; not even the best base model is able to match any of the reasoning LLMs (cf. Tab. 4). This, however, comes at a steep cost as moving from GPT-4o to o3 doubles accuracy, but considerably increases the number of completion tokens (1777%) and thus the computational costs (1034%). Moreover, scaling test-time compute on the same model also boosts overall performance, but does not guarantee higher accuracy across all tasks. For example, while o4-mini-high typically outperforms o4-mini (LRL: 12.8 vs. 12.3), it underperforms on the medium complexity tier (40% vs. 52%). This plateau effect demonstrates that, beyond a certain threshold, additional compute may yield diminishing or even negative returns.
⟦!t⟧ Refer to captionwidth=⟧figs/completion_tokens_vs_accuracy_combined.pdf
Scaling Model Parameters Brings Limited Gains in Logical Reasoning.
Increasing model size yields only marginal gains in logical reasoning. While larger models like Llama-3-70B and GPT-4.5-prev generally outperform their smaller counterparts, returns diminish as improvements are increasingly modest (see Tab. 4). As even the largest base models still fall short of the reasoning models, a capability gap remains that suggests that scaling model parameters alone does not guarantee substantial advances in logical reasoning capabilities.
Compute Increases with Task Complexity.
As reasoning tasks become more complex, reasoning LLMs require more tokens to solve the tasks, leading to increased financial costs (see Fig. 5). These increased demands impose limits on scaling reasoning through increased test-time compute.
6 Conclusion and Future Direction
Our evaluations reveal that while current LLMs readily produce syntactically valid logic rules, robust logical reasoning remains elusive for conventional LLMs, especially as task complexity scales. Scaling model parameters yields only limited gains. Reasoning LLMs, aided by increased test-time compute, close part of this gap, albeit at significant computational costs.
Notably, curriculum learning on SLR-Bench significantly boosts both in-domain and downstream reasoning. Our SLR-tuned Llama3-8B not only outperforms all conventional LLMs on SLR-Bench, but also surpasses several SOTA reasoning LLMs at a fraction of their inference costs. Furthermore, we observe improved reasoning capabilities across a wide range of established benchmarks, underscoring the effectiveness of curriculum-based logic-tuning for downstream reasoning tasks.
Looking ahead, SLR paves the way for several promising research directions, including the integration of reinforcement learning for reasoning LLMs using SLR, expanding into richer logical domains, benchmarking neuro-symbolic and interactive reasoning systems, and moving toward higher-order logic tasks like causal inference. Ultimately, SLR provides a flexible and extensible resource for probing and advancing the frontiers of logical reasoning in the next generation of LLMs.
7 Limitations
Acknowledgements
We acknowledge support of the hessian.AI Innovation Lab (funded by the Hessian Ministry for Digital Strategy and Innovation), the hessian.AISC Service Center (funded by the Federal Ministry of Education and Research, BMBF, grant No 01IS22091), and the Centre for European Research in Trusted AI (CERTAIN). Further, this work benefited from the ICT-48 Network of AI Research Excellence Center “TAILOR” (EU Horizon 2020, GA No 952215), the Hessian research priority program LOEWE within the project WhiteBox ⟦GA No LOEWE/ 2/13/519/03/06.001(0010)/77⟧, the HMWK cluster projects “Adaptive Mind” and “Third Wave of AI”, and from the NHR4CES. This work was supported by the Priority Program (SPP) 2422 in the subproject “Optimization of active surface design of high-speed progressive tools using machine and deep learning algorithms“ funded by the German Research Foundation (DFG). Further, this work was funded by the European Union (Grant Agreement no. 101120763 - TANGO) as well as the AlephAlpha Collaboration lab 1141. Views and opinions expressed are, however, those of the author(s) only and do not necessarily reflect those of the European Union or the European Health and Digital Executive Agency (HaDEA). Neither the European Union nor the granting authority can be held responsible for them.
Broader Impact
Appendix A Task Specification

### v5

@@ 4 SLR-Bench: Instantiating SLR
Design rationale. The logic task is inspired by the V-LoL trains domain helff2025vlol; michalski1980; Mitchell1997MachineLI, chosen for three main reasons. First, its hierarchical structure (trains $\rightarrow$ cars $\rightarrow$ attributes) naturally yields first-order rules richer than simple lookups yet more tractable than full theorem proving. Second, its small, discrete attribute domains enable precise control over task complexity. Third, it offers a clean proof of concept for scalable logic synthesis, being fully synthetic and easily extendable to new domains.
⟦t⟧ Curriculum Learning (SLR-Bench) LRL ($\uparrow$0-20) Syntax($\uparrow$%) Basic($\uparrow$%) Easy($\uparrow$%) Medium($\uparrow$%) Hard($\uparrow$%) Llama3.1-8b-it 3.8 80 70 7 0 0 Llama3.1-8b-it-SLR 8.7 (+4.9) 100 (+20) 96 (+26) 54 (+47) 18 (+18) 5 (+5) Downstream Reasoning Performance SLR-Homes ($\uparrow$%) MMLU ($\uparrow$%) MMLU-Stats ($\uparrow$%) CLUTRR ($\uparrow$%) LogiQA ($\uparrow$%) LogiQA2 ($\uparrow$%) Llama3.1-8b-it 10.3 63.3 42.6 25.5 30.1 34.3 Llama3.1-8b-it-SLR 42.3 (+32) 66.1 (+2.8) 59.7 (+17) 32.1 (+6.6) 31.0 (+0.9) 39.4 (+5.2) GPQA ($\uparrow$%) GPQA-Ext. ($\uparrow$%) GPQA-Dia. ($\uparrow$%) ARC-Easy ($\uparrow$%) ARC ($\uparrow$%) HellaSwag ($\uparrow$%) Llama3.1-8b-it 31.7 26.9 21.7 81.4 52.7 57.4 Llama3.1-8b-it-SLR 32.8 (+1.1) 33.0 (+6.1) 28.3 (+6.6) 82.8 (+1.4) 54.6 (+1.9) 58.9 (+1.5)
Table 2: Curriculum Learning and Generalization. Benchmark scores ($\uparrow$) for base and SLR-tuned models on SLR-Bench and downstream benchmarks; LRL measuring cumulative curriculum progress. The tuned model surpasses the baseline across all curriculum stages, while generalizing to other downstream reasoning tasks.
Languages. Each curriculum level is parameterized by level-specific language $\mathcal{L}$, detailed in App. B.2. The vocabulary includes mutually exclusive class labels eastbound and westbound, which serve as the targets for classification tasks $(E^{+},E^{-})$. Background knowledge $B$ is sampled from a vocabulary of five predicates (has_car, car_num, car_color, car_len, and has_wall) with their respective grounding domains defined in App. B.2. As curriculum levels increase, the vocabulary expands monotonically by introducing new predicates and grounding domains selected from a predefined set. Semantic coherence is ensured by constraining predicate groundings (e.g., only colors as arguments for car_color), and by enforcing mutually exclusive constraints across predicates (e.g., passenger cars cannot carry payloads).
Task configs. Each curriculum level is parameterized by level-specific settings of $\theta$, summarized in App. Tab. 4 and supplied directly to the synthesizer (Alg. 1). Problem size ($\kappa$) increases steadily across levels, maintaining an equal balance of positive and negative samples. Levels 1–5 use a mirror sampling policy for background knowledge, generating simple, nearly identical east- and westbound trains that differ only in ground atoms relevant to $R^{\star}$. From level 6 onward, the background is sampled uniformly from the filtered Herbrand base, increasing diversity. Rule generation is uniform for the basic levels; from level 6, 30% of rules are LLM-guided, introducing greater variety in variables, arithmetic, recursion, and more.
SLR-Curriculum. SLR-Bench comprises 20 levels across four tiers: basic, easy, medium, and hard. Each level systematically increases complexity by expanding task size ($\kappa$), adding new car constants and predicates, lengthening rules, and varying both the background knowledge and rule sampling policy; see App. Sec. B, Tab. 4. As a result, the combinatorial space of possible tasks grows exponentially, and later levels become progressively harder and require deeper reasoning beyond surface cues.
SLR-Homes. Alongside the main curriculum, we introduce an OOD dataset that instantiates SLR in a new symbolic domain. It uses a distinct vocabulary, relational structure, and target concept to define house classification tasks based on attributes. SLR-Homes enables the assessment of cross-domain generalization and comprises 500 tasks across five difficulty levels, increasing in complexity.
Intended Use. SLR-Bench is designed for two complementary purposes. (1) As a static benchmark, it enables fine-grained evaluation of an LLM’s reasoning abilities across tasks of increasing logical complexity. It is also easily extensible to accommodate future improvements in model capabilities. (2) As a dynamic curriculum, it serves as a training backbone, supplying structured reasoning tasks and feedback to enhance reasoning in both conventional and reasoning LLMs.
5 LLMs Can’t Do Induction at Scale
We benchmark and train LLMs on SLR-Bench, assessing reasoning, syntactic correctness, and computational efficiency across the four difficulty tiers: basic, easy, medium, and hard. Our analysis highlights key trends, common failure modes, and the effectiveness of curriculum-based logic-tuning.
Training Setup. We investigate how LLMs benefit from curriculum training on SLR-Bench with SFT (for more details see App. Sec. D). To prevent data leakage, we ensure that no prompts or rules from the test set are included in the training set.
Evaluation Setup. All models are evaluated in a zero-shot setting using SLR-Bench prompts, with a single attempt per task (pass@1). We report the following metrics: (i) Logical Reasoning Level (LRL): A cumulative score over all curriculum levels $L$, where $\text{solved}_{\ell}$ and $\text{tasks}_{\ell}$ denote the number of solved and total tasks at level $\ell$: $\text{LRL}=\sum_{\ell=1}^{L}\frac{\#\text{solved}_{\ell}}{\#\text{tasks}_{\ell}}$ (ii) Syntax Score: The proportion of predicted logic rules that are syntactically valid. (iii) Logical-Reasoning Accuracy: The fraction of correct solutions per complexity tier. (iv) Compute: The aggregate completion tokens and computational cost for each the models. For further details on downstream evaluations, see App. Sec. D, and for pricing, refer to App. Tab. 5.
5.1 Analysis and Key Findings
[…]
In this section, we analyze the key effects of SLR training. We first report downstream reasoning gains for models trained with SLR, namely Llama-3.1-8B-IT (Tab. 2) and Qwen3-8B (App. Tab. 10). We then benchmark SOTA conventional and reasoning LLMs on SLR-Bench (Tab. 3), with an extended leaderboard provided in App. Tab. 16.
SLR Boosts Downstream Reasoning. Curriculum learning on SLR-Bench yields substantial gains in both in-domain and downstream reasoning (Tab. 2). SLR-tuned models consistently outperform conventional LLMs on SLR-Bench (cf. Tab. 3) and even surpass dedicated reasoning models such as Gemini-2.0-flash-thinking, while requiring significantly fewer inference tokens and lower compute. Performance on SLR-Homes further demonstrates robust zero-shot transfer to previously unseen domains, suggesting that SLR learns domain-agnostic reasoning capabilities rather than dataset-specific heuristics. Beyond SLR-Bench, these gains transfer to a broad set of established reasoning benchmarks, including MMLU High School Statistics (+17) hendryckstest2021, CLUTRR sinha2019clutrr, LogicQA liu2020logiqa, LogicQA2 Hanmeng2023logicqa2, ARC Clark2018ThinkYH, HellaSwag zellers2019hellaswag, and GPQA variants rein2024gpqa. We hypothesize that SLR training introduces an inductive bias toward systematic rule induction over structured relational contexts. Models are encouraged to internalize abstractions for identifying relational structure, composing and applying reasoning patterns. These abstractions generalize to downstream deductive and knowledge-intensive tasks, where performance likewise depends on structured reasoning.
Curriculum Order Matters. An ablation study (App. B.3) examines how curriculum order affects training performance by comparing ordered, random, and reverse progressions. Training with the ordered curriculum achieves the best results, confirming that structured progression supports smoother and more stable learning.
Curriculum Levels Modulate Task Complexity: LLMs Break Down as Complexity Increases. SLR-Bench creates a controlled gradient in logical complexity as model performance steadily declines throughout the curriculum levels (cf. Fig. 3, Tab. 3). Most models readily solve basic levels. Base LLMs already falter on easy ones, solving fewer than half. Reasoning LLMs perform better, yet their accuracy drops sharply at the medium tier, and none succeed on hard. This pattern is also reflected in the LRL score, empirically indicating how far each model can progress before performance collapses. An ablation study (App. E.1) further shows how increasing rule length or problem size, and switching from “uniform” to “LLM-guided” or “mirror” to “uniform” sampling, influences task complexity, confirming that SLR provides fine-grained control over logical complexity.
Test-time Scaling Improves Reasoning, but Returns Diminish and Costs Escalate. Reasoning LLMs clearly outperform the base models; not even the best base model is able to match any of the reasoning LLMs (cf. Tab. 3). CoT prompting provides modest gains ($\approx 5\%$, App. F.4). Further, scaling test-time computation comes at a steep cost as moving from GPT-4o to o3 doubles accuracy, but considerably increases the number of completion tokens (1777%) and thus the computational costs (1034%). Moreover, scaling test-time compute on the same model also boosts overall performance (see App. E.3), but does not guarantee higher accuracy across all tasks. E.g., while o4-mini-high typically outperforms o4-mini (LRL: 12.8 vs. 12.3), it underperforms on the medium tier (40% vs. 52%). This plateau effect demonstrates that, beyond a certain threshold, additional compute may yield diminishing or even negative returns.
Scaling Model Parameters Brings Limited Gains in Logical Reasoning. Increasing model size yields only marginal gains in logical reasoning. While larger models like Llama-3-70B and GPT-4.5-prev generally outperform their smaller counterparts, returns diminish as improvements are increasingly modest (see Tab. 3). As even the largest base models still fall short of the reasoning models, a capability gap remains that suggests that scaling model parameters alone does not guarantee substantial advances in logical reasoning capabilities.
Common Error Modes. Base LLMs reliably produce syntactically valid rules, as reflected by their high syntax scores (Tab. 3). Reasoning models achieve slightly lower syntax scores, particularly on higher-complexity tasks, where longer generations increase the likelihood of invalid outputs. However, the dominant performance bottleneck is semantic rather than syntactic, as evidenced by the persistent gap between syntax and LRL. A detailed error analysis (App. G) identifies five recurring failure modes: (i) over-generalized rules, (ii) under-generalized rules, (iii) failure of abstraction, (iv) degenerate reasoning loops, and (v) shortcut behavior in SOTA reasoning models (GPT-5, and GPT-5-mini). Additionally, accuracy decreases with increasing output length (App. Tab. 15), suggesting that extended reasoning traces often lead to derailment instead of deeper reasoning.
6 Conclusion and Future Direction
Our evaluations reveal that while current LLMs readily produce syntactically valid logic rules, robust logical reasoning remains elusive for conventional LLMs, especially as task complexity scales. Scaling parameters yields limited gains. Reasoning LLMs, aided by increased test-time compute, close part of this gap, albeit at significant computational costs. Notably, curriculum learning on SLR-Bench substantially boosts in-domain and downstream reasoning. Specifically, our SLR-tuned Llama3-8B outperforms most conventional LLMs on SLR-Bench and surpasses several SOTA reasoning LLMs at a fraction of their inference costs.
Looking ahead, SLR enables diverse extensions, including the integration of reinforcement learning, richer logical domains, benchmarking neuro-symbolic and interactive reasoning systems, and advancing toward higher-order tasks such as causal inference. While SLR-Bench intentionally focuses on clean, single-rule ILP tasks with a perfect symbolic judge to isolate reasoning ability, future work can incorporate controlled forms of real-world noise, ambiguity, and multi-rule interactions to systematically study robustness alongside reasoning capabilities. Overall, SLR provides a flexible and extensible testbed for probing and improving LLM reasoning.
7 Limitations
Acknowledgments
We acknowledge support of the hessian.AI Innovation Lab (funded by the Federal Ministry of Research, Technology and Space, BMFTR, grant no. 16IS22091), the hessian.AISC Service Center (funded by the Federal Ministry of Education and Research, BMBF, grant No 01IS22091), and the Center for European Research in Trusted AI (CERTAIN). Further, this work benefited from the ICT-48 Network of AI Research Excellence Center “TAILOR” (EU Horizon 2020, GA No 952215), the Hessian research priority program LOEWE within the project “WhiteBox”, the HMWK cluster projects “Adaptive Mind” and “Third Wave of AI”, and from the NHR4CES. This work has also benefited from the BMWE project "EU-SAI: Souveräne KI für Europa," 13IPC040G, and also from early stages of the Cluster of Excellence "Reasonable AI" funded by the German Research Foundation (DFG) under Germany’s Excellence Strategy— EXC-3057; funding will begin in 2026. This work was supported by the Priority Program (SPP) 2422 in the subproject “Optimization of active surface design of high-speed progressive tools using machine and deep learning algorithms“ funded by the German Research Foundation (DFG). Further, this work was funded by the European Union (Grant Agreement no. 101120763 - TANGO) as well as the AlephAlpha Collaboration lab 1141. This work was supported in part by OpenAI Research Credits.
Broader Impact
@@ Acknowledgments
Appendix A Extended Notation
Appendix B SLR-Bench Details
B.1 Task Specification